OpenStudy (anonymous):

Complex number Help! Find the real and imaginary parts of the following numbers and plot them in the complex plane. z=e^(2-pi*i/4)

OpenStudy (freckles):

\[z=e^{2-\frac{\pi}{4} i} =e^2 e^{\frac{-\pi}{4}i} \text{ by law of exponents } \]

OpenStudy (anonymous):

I had that so far, but I don't know where to go from there.

OpenStudy (freckles):

\[e^{ \theta i }=\cos(\theta)+i \sin(\theta) \\ \text{ so } a \cdot e^{i \theta} =a \cos(\theta)+ i a \sin(\theta)\]

OpenStudy (freckles):

real part is a cos(theta) imaginary part is a sin(theta)

OpenStudy (freckles):

cool or not cool?

OpenStudy (anonymous):

So would the real part be e^(2)cos(theta)?

OpenStudy (anonymous):

and the imaginary part be -pi/4sin(theta)?

OpenStudy (freckles):

I think you are working too fast

OpenStudy (freckles):

your theta here is definitely -pi/4 and your a is what was in front of the e^(i theta) thing

OpenStudy (anonymous):

I don't really understand then.

OpenStudy (anonymous):

e^2 then?

OpenStudy (anonymous):

I mean e^2 is a.

OpenStudy (freckles):

e^2 is a -pi/4 is theta

OpenStudy (anonymous):

so the real part is e^2*cos(-pi/4)?

OpenStudy (freckles):

yep

OpenStudy (freckles):

and you can evaluate the cos(-pi/4) part

OpenStudy (freckles):

\[e^2 \frac{\sqrt{2}}{2} \text{ is real part }\]

OpenStudy (anonymous):

oh okay. so the imaginary would be e^2*sin(-pi/4)?

OpenStudy (anonymous):

and you would evaulate that too.

OpenStudy (freckles):

so in general assume x and y are real \[e^{x+yi}=e^{x}e^{yi}=e^x (\cos(y)+i \sin(y))= e^x \cos(y)+i e^x \sin(y) \\ \text{ where the real part is } e^x \cos(y) \\ \text{ and the imgainry part is } e^x \sin(y)\]

OpenStudy (freckles):

and yes to your question you can also evaluate that

OpenStudy (anonymous):

Okay. That makes a lot more sense! Thank you!

OpenStudy (anonymous):

Can I ask you about the next one I have if you're still there?

OpenStudy (freckles):

yes i was gone but not i'm here

OpenStudy (freckles):

now*

OpenStudy (anonymous):

The next one is \[z=e^{-1-i}\] I set it up like the last one where \[z=\frac{ e^{-1} }{ e^{i}}\]

OpenStudy (freckles):

hey are we doing the same thing?

OpenStudy (anonymous):

yeah, sorry I didn't say that.

OpenStudy (anonymous):

The thing that confuses me is there is no theta

OpenStudy (anonymous):

Unless 1 is theta

OpenStudy (freckles):

you can do the exact same I did above

OpenStudy (freckles):

\[e^{-1-i}=e^{-1}e^{-i}\]

OpenStudy (freckles):

your theta is -1 in this case

OpenStudy (freckles):

\[e^{-1}e^{-1 \cdot i}\]

OpenStudy (anonymous):

Oh okay. I can do it from there. thanks!